# Rethink the Evaluation Protocol of Model Merging on Classification Task

### v1

@@ 3.3 Alignment Formulation
+FT-C 87.5 58.2 52.2 69.1 97.7 73.0
Table 1: Comparison of FT-Classifier evaluation with the corresponding current one. The results in this Table are based on a few-shot setting of $k=5$. FT represents the results of the Fine-tuned model, MTL represents the results of traditional Multi-task Learning, WA represents Weight Averaging, TA represents Task Arithmetic, and +FT-C represents +FT-Classier evaluation protocol.
@@ 3.4 Further Discussion
Model merging involves combining the parameters of multiple $f_{t}$ models to generate $f_{m}$. Consequently, the output of $f_{m}$ is inevitably shifted in the latent space compared to the original $f_{t}$. This shift is unavoidable. In this case, using $\mathbf{W}_{t}$, which is aligned well with $f_{t}$, to evaluate $f_{m}$ is inappropriate. The evaluation result reflects not the capability of $f_{m}$, but rather its compatibility with the fine-tuned classifier $\mathbf{W}_{t}$. Moreover, through experiments, we find that simply aligning the merging outputs and the classifier in the latent space can significantly improve the classification performance of $f_{m}$. Naturally, this motivates us to propose a new and reasonable evaluation protocol for model merging. This protocol would not only enable a more precise assessment of $f_{m}$, but also enhance classification performance through mitigating the misalignment issue.
4 FT-Classifier Evaluation Protocol for Alleviating the Misalignment Issue
In model merging, obtaining few-shot data for each task is a relatively flexible requirement. Many studies use few-shot data to guide the merging methods Xiao et al. (2023), and some even rely on large-scale labeled or unlabeled data to determine merging weights Yang et al. (2024c); Jin et al. (2022) or the parameters to be pruned Ortiz-Jimenez et al. (2024); Zhu et al. (2024). In our proposed evaluation protocols, we also incorporate few-shot samples for misalignment alleviation. Meanwhile, as the goal of model merging is to produce a single model capable of flexibly addressing various tasks, we argue that the evaluation protocol should avoid introducing additional parameters. Hence, we select the Aligned $\mathbf{W_{t}^{\prime}}$ variants to mitigate the misalignment.
In particular, we adopt few-shot samples to fine-tune an aligned classifier for merging model evaluation. Compared to the current evaluation protocol, this approach introduces an additional post-processing step, while keeping the parameter scale and network structure unchanged. Since only a few-shot samples are used for training, the required time remains entirely acceptable. The training objective of this approach is consistent with Eq. 8, also called FT-classifier Evaluation, with the only difference being the use of few-shot data.
Comparable results of the current evaluation and FT-Classifier are presented in Figure 5 and Table 1. From the Figure, we observe that after simple fine-tuning even on 1-shot samples, FT-Classifier Eval performs significantly better than the Current Eval. This result leads us to wonder whether the alignment is truly straightforward; otherwise, it should not have been captured so easily.
Refer to caption Refer to caption Refer to caption
@@ 4 FT-Classifier Evaluation Protocol for Alleviating the Misalignment Issue
Figure 5: The performance of different evaluation protocols under varying numbers of few-shot examples $k$. Corresponding detailed results while $k=5$ are shown in Table 1.
5 The Misalignment is an Orthogonal Transformation
Intuitively, we hypothesize that this transformation is the simplest form of an orthogonal matrix, meaning that the merging outputs can be aligned in the latent space through simple rotations or reflections. To accomplish this, we enforce $\mathbf{M}_{t}$ to be an orthogonal matrix by adding a regularization constraint while training. Specifically, we introduce a regularization loss on Aligned $\mathbf{M}_{t}$ to ensure this property. The objective function is defined as follows:
[…]
where $\alpha$ is the weight hyperparameter, and $\mathbf{I}$ is an identity matrix.
Method Unaligned + Orth $\mathbf{M}_{t}$ + $\mathbf{M}_{t}$
WA 55.8 72.8 74.9
TA 64.3 72.7 74.5
Ties 65.1 73.8 74.7
Table 2: Comparison of realigned current eval, orthogonal transformation alignment (+Orth $\mathbf{M}_{t}$), and normal alignment (+ $\mathbf{M}_{t}$).
Detailed results of the orthogonal regularization are shown in Table 2. From the Table, we observe that when $\mathbf{M}_{t}$ converges to an orthogonal transformation, the final classification performance is not significantly affected (only nearly 2% dropped). This demonstrates that through an orthogonal transformation, the merging outputs can be aligned with the fine-tuned classifier. Since orthogonal transformations in space are rigid transformations, such as rotations or reflections, this suggests that these merging methods preserve the essential capabilities of the fine-tuned models, while the intrinsic properties of the manifold in the latent space remain unchanged.
@@ 6 Experiments
The above observations demonstrate that alleviating the misalignment and modifying the evaluation protocol is necessary. And our new evaluation protocols more accurately assess the performance of merging methods and are able to improve the overall classification performance. In this Section, we will provide more detailed analyses through additional supplementary experiments.
6.1 Comparison on NLP Tasks
6.1.1 Datasets
@@ 6.1.2 Implementation Details
For the backbone of models, $f_{b}$, we primarily use bert-base-uncased Devlin et al. (2018) as the base architecture, and the fine-tuned checkpoints are referenced from Yang et al. (2024b). Additionally, for a more comprehensive evaluation, we also adopt T5-base and T5-large Raffel et al. (2020) as $f_{b}$. For the scaling hyperparameters of the Task Arithmetic and Ties models, we follow the original paper by using a portion of labeled data for selection. These parameters are then fixed across all evaluation protocols to ensure a fair comparison. Our code is released in Github  [^1] https://github.com/fskong/FT-Classifier-for-Model-Merging.
6.1.3 Results Analysis
@@
Each merging model is evaluated on all datasets using two protocols: Current Eval and FT-Classifier Eval, to calculate accuracy. To save space, we compute the average accuracy and present it using a line chart, with detailed results provided in the Appendix. To demonstrate that the FT-Classifier Eval maintains a certain level of robustness even with a small number of few-shot samples, we conduct experiments using five random seeds and report the mean and variance in the Figure.
@@
The experimental results using BERT as $f_{b}$ have been reported in Figure 5 and Table 1. Unless otherwise specified, the experiments in this paper are based on BERT. And the results conducted on T5-base and T5-large are shown in Figure 6. From the Figure, we find that the main conclusions are essentially consistent with those on BERT, strengthening our arguments.
Refer to caption Refer to caption Refer to caption
(a) NLP tasks performance on T5-base.
Refer to caption Refer to caption Refer to caption
(b) NLP tasks performance on T5-large.
Figure 6: The performance of different evaluation protocols on NLP tasks under T5-base and T5-large. Detailed results are reported in Appendix A.3.
Refer to caption Refer to caption Refer to caption
(a) CV tasks performance on ViT-B-16.
Refer to caption Refer to caption Refer to caption
(b) CV tasks performance on ViT-B-32.
Figure 7: The performance of different evaluation protocols on CV tasks under ViT-B-16 and ViT-B-32. Detailed results are reported in Appendix A.4.
6.2 Comparison on CV Tasks
6.2.1 Datasets
6.2.2 Implementation Details
For $f_{b}$, we primarily utilize ViT-B-16 and ViT-B-32 Radford et al. (2021) as the base architecture, and the fine-tuned checkpoints are referenced from Yang et al. (2024b). The selection for the merging scale is consistent with the description in the NLP tasks. It is worth noting that we found that, possibly due to the large number of classes in most CV datasets, using few-shot samples is insufficient to fine-tune the classifier effectively. Therefore, we increased the number of samples used for alignment learning. Specifically, we sampled subsets from the test set with sampling rates of {0.5%, 1%, 2.5%, 5%, 10%}.
6.2.3 Results Analysis
To verify that the misalignment issue also exists in CV tasks, we visualize the image embeddings using T-SNE. The results are shown in Appendix A.2. From the Appendix, we observe that the merging outputs also exhibit good clustering, indicating their high quality. Therefore, we believe that the misalignment problem also exists in CV tasks.
Further, to mitigate the misalignment issue in CV tasks, we applied the FT-Classifier protocol to re-evaluate different merging methods. The comparable results are shown in Figure 7. We find that with just over 1% of unlabeled data, the misalignment problem can be alleviated. This not only demonstrates the effectiveness of the FT-Classifier but also implicitly proves that misalignment is easy to capture.
6.3 Difference between $f_{m}$ and $f_{b}$
@@
Previous analyses demonstrate that different $f_{m}$ achieves significant improvements when evaluated with a simple fine-tuning classifier. To further verify that the performance improvement stems from the inherent capability of the $f_{m}$ itself rather than additional information introduced by fine-tuning, we tested the results of the original pre-trained models $f_{b}$ under different evaluation protocols. The experimental results are presented in Table 3.
@@
From the Table, we observe that regardless of the evaluation protocol, $f_{m}$ consistently outperforms $f_{b}$. This demonstrates that all $f_{m}$ models effectively incorporate knowledge from fine-tuned models. Meanwhile, our new evaluation protocols do not show significant improvements over the original protocol when applied to $f_{b}$, indicating that models with limited parameters cannot retain sufficient task-specific information. This proves that our evaluation protocols are essentially fair compared to the original ones and just activate the fine-tuned information that $f_{m}$ already contains.
@@ 7 Related Works
The task vector indicates the offset of a certain task based on the pre-trained model in the model parameter space. We introduce the following latest methods except for the representative referred to before. Motivated by model pruning that eliminating unnecessary parameters could also maintain performance, Dare Yu et al. (2024) reduces potential conflicts in the delta parameters by randomly dropping. To mitigate forgetting while training, MagMax Marczak et al. (2024) combines sequential fine-tuning with maximum-magnitude weight selection to enable effective knowledge integration across tasks. TALL-masks Wang et al. (2024) localizes task-specific information within the multi-task vector by deactivating irrelevant parts for each task using binary masks. Model Tailor Zhu et al. (2024) derives a sparse mask to identify the "model patch" using a fusion of salience and sensitivity analysis, then introduces a compensation mechanism to "decorate the patch", enhancing performance on both target and original tasks.
@@ 8 Conclusion
Model merging combines multiple fine-tuned models into a single one in the parameter space, enabling multi-task ability without full datasets or extra fine-tuning. We identify the misalignment issue between merging outputs and fine-tuned classifiers in classification tasks and show that addressing it improves classification performance. We also enhance evaluation protocols by fine-tuning an aligned classifier with few-shot examples, boosting evaluation quality with minimal time cost. Finally, we prove this misalignment can be converged as an orthogonal transformation, offering insights into model merging’s intrinsic properties.
@@ A.1 Visualization of Sentence Embedding
A.2 Visualization of Image Embedding
The image embeddings produced by the $f_{m}$ model are shown in Figure 9 and 10. It is worth noting that, due to the large number of classes in the CV datasets, we did not display the T-SNE results of the fine-tuned model for better visualization. As shown in the Figure, we find that the embeddings produced by different $f_{m}$ outputs also exhibit high quality, which is consistent with the conclusions drawn from the NLP tasks.
@@ A.2 Visualization of Image Embedding
A.3 Detailed Results on NLP Tasks
Detailed results on NLP tasks by T5-base and T5-large are presented in Table 4.
(a) T5-base
Method AG News Yelp Amazon Yahoo DBpedia Average
Fine-tuned 94.5 66.6 62.9 75.2 99.1 79.6
Weight Averaging 79.0 51.8 51.7 42.2 79.9 60.9
+FT-Classier 85.3 60.5 55.3 61.8 97.1 72.0
Task Arithmetic 80.6 59.2 55.2 51.5 79.5 65.2
+FT-Classier 88.2 64.0 58.2 67.6 98.4 75.3
Ties 80.0 59.2 55.4 50.7 79.3 64.9
+FT-Classier 88.3 63.8 58.1 68.0 98.4 75.4
(b) T5-large
Method AG News Yelp Amazon Yahoo DBpedia Average
Fine-tuned 94.9 68.6 64.9 77.1 99.1 80.9
Weight Averaging 79.2 55.5 44.3 49.0 94.5 64.5
+FT-Classier 88.0 62.6 59.7 67.9 98.2 75.3
Task Arithmetic 90.8 58.3 60.5 72.9 92.7 75.1
+FT-Classier 91.2 65.9 61.9 73.5 98.7 78.3
Ties 91.3 56.0 60.3 73.2 90.4 74.3
+FT-Classier 88.4 63.9 58.2 68.0 98.4 75.4
Table 4: Comparison on NLP tasks by T5. The results in this Table are based on a few-shot setting of $k=20$.
A.4 Detailed Results on CV Tasks
Detailed results on CV tasks by ViT-B-16 and ViT-B-32 are presented in Table 5.
(a) ViT-B-16
Method SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Average
Fine-tuned 81.8 86.8 96.9 99.7 97.8 99.1 99.7 82.0 92.9
Weight Averaging 67.7 70.0 75.3 79.5 74.9 60.1 94.4 43.8 70.7
+FT-Classier 69.7 72.7 83.8 89.7 73.7 84.3 95.3 51.5 77.6
Task Arithmetic 61.1 65.9 74.0 76.2 88.0 73.9 98.4 53.0 73.8
+FT-Classier 68.1 72.3 86.3 90.8 90.2 90.3 98.8 64.2 82.7
Ties 69.1 72.5 80.5 84.0 85.0 71.5 98.1 54.9 77.0
+FT-Classier 71.3 74.1 86.0 90.6 90.4 90.2 98.8 64.5 83.2
(b) ViT-B-32
Method SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Average
Fine-tuned 75.3 77.7 96.1 99.7 97.5 98.7 99.7 79.4 90.5
Weight Averaging 65.3 63.4 71.4 71.7 64.2 52.8 87.5 50.1 65.8
+FT-Classier 71.5 73.6 86.7 91.9 79.5 83.1 97.1 58.7 80.3
Task Arithmetic 55.2 54.9 66.7 78.9 80.2 69.7 97.3 50.4 69.1
+FT-Classier 62.5 61.5 80.4 89.9 86.4 89.9 98.4 61.3 78.8
Ties 65.0 64.4 74.8 77.4 81.2 69.3 96.5 54.5 72.9
+FT-Classier 73.1 77.0 89.1 93.9 88.0 89.7 98.7 68.1 84.7
Table 5: Comparison on CV tasks by ViT. The results in this Table are based on a sampling rate of 10%.

### v2

@@ 4.1 Different Transformation Hypothesis
+ KNN Eval 87.0 $\pm$ 1.6 54.6 $\pm$ 2.9 49.5 $\pm$ 1.5 63.4 $\pm$ 0.6 97.1 $\pm$ 0.2 70.3 $\pm$ 1.2
+ FT-Classifier Eval 88.1 $\pm$ 0.5 57.1 $\pm$ 0.9 52.6 $\pm$ 0.6 69.1 $\pm$ 0.6 97.7 $\pm$ 0.1 72.9 $\pm$ 0.1
Table 2: Comparison of FT-Classifier Eval and KNN Eval with the corresponding Current Eval (represented by merging method name). The results in this Table are based on a few-shot setting of $k=5$.
@@ 4.2 Further Discussion
Since orthogonal transformations in space are rigid, such as rotations or reflections, these transformations do not alter the intrinsic structure or geometric properties of the manifold. This demonstrates that the merging outputs largely retain the classification characteristics of the fine-tuned outputs. By further analyzing Figure 4 and considering the properties of orthogonal transformations, we can draw the following conclusions:
(1) With only a simple transformation with a limited number of parameters, the misalignment can be captured, achieving surprising classification performance that approaches the upper bound of the fine-tuned model. This further explains why merging outputs deliver strong performance even under direct classification (e.g., KNN Eval), as they preserve the essential information necessary for effective classification.
(2) The outputs of different $f_{m}$ contain comparable classification information, and Current Eval does not effectively assess $f_{m}$. Since the performance gap of different Aligned $f_{m}$ under the same transformation is much smaller than $f_{m}$. For example, the performance of Weight Averaging was initially around 10% lower than that of Task Arithmetic, but after alignment, the performance of Aligned Weight Averaging and Aligned Task Arithmetic becomes nearly identical.
These findings demonstrate that the misalignment issue hinders the Current Eval from effectively assessing the capacity of $f_{m}$. Clearly, alleviating the misalignment problem could potentially improve classification performance. Therefore, we hope to explore a new evaluation protocol that can effectively assess different merging methods by alleviating the misalignment.
5 FT-Classifier Evaluation Protocol
@@
Figure 5: The performance of FT-Classifier Eval under varying numbers of few-shot examples $k$. Corresponding detailed results while $k=5$ are shown in Table 2.
Model merging combines the parameters of multiple fine-tuned models $f_{t}$ to produce a merged model $f_{m}$. Consequently, as discussed before, the output of $f_{m}$ inevitably shifts in the latent space relative to the original $f_{t}$, making this shift unavoidable. In this case, using $\mathbf{W}_{t}$, which is aligned well with $f_{t}$, to evaluate $f_{m}$ is inappropriate. The evaluation result reflects not the capability of $f_{m}$, but rather its compatibility with the fine-tuned classifier $\mathbf{W}_{t}$. Moreover, through experiments, we find that simply aligning the merging outputs and the fine-tuned outputs in the latent space can significantly improve the classification performance of $f_{m}$. Naturally, this motivates us to propose a new and reasonable evaluation protocol, which addresses the misalignment issue to more accurately assess $f_{m}$ and achieve improved performance.
Revisiting the objective of model merging, which aims to generate a new model $f_{m}$ using $f_{1},...,f_{T}$ that performs well across all $T$ tasks, we note that the upper bound of $f_{m}$’s performance on a specific task $t$ corresponds to the result from fine-tuned model $f_{t}$. Therefore, the objective of model merging is to minimize this performance discrepancy of $f_{m}$ and $f_{t}$ Ortiz-Jimenez et al. (2023). By optimizing this discrepancy, we could directly train an aligned classifier for $f_{m}$. That is, in the new evaluation protocol, we introduce an additional post-processing step while keeping the parameter scale and network structure unchanged, referred to as FT-Classifier Eval, which is defined as:
$\arg\min_{\mathbf{W}_{t}^{\prime}}\sum_{t=1}^{T}g(f_{m}(X_{t};\mathbf{\Theta}_{m})\!^{\top}\!\mathbf{W}_{t}^{\prime},f_{t}(X_{t};\mathbf{\Theta}_{t})\!^{\top}\!\mathbf{W}_{t})$ (8)
where $g$ is a measurement function. In this objective, we instantiate $g$ as KL divergence Kullback and Leibler (1951), and the new classifier $\mathbf{W}_{t}^{\prime}$ is initialized with the value of $\mathbf{W}_{t}$.
Refer to caption Refer to caption
(a) Performance on T5-base (b) Performance on T5-large
Figure 6: The performance of different evaluation protocols under T5-base and T5-large.
In model merging, obtaining few-shot data for each task is a relatively flexible requirement. Many studies use few-shot data to guide the merging methods Xiao et al. (2023), and some even rely on large-scale labeled or unlabeled data to determine merging weights Yang et al. (2024c); Jin et al. (2022) or the parameters to be pruned Ortiz-Jimenez et al. (2023); Zhu et al. (2024). Therefore, we also incorporate few-shot samples for aligned classifier training in FT-Classifier Eval.
Comparable results of Current Eval and FT-Classifier Eval are presented in Figure 5, and detailed results, including KNN Eval, utilizing 5-shot samples are shown Table 2. From the Figure, we observe that after simple fine-tuning even on 1-shot samples, FT-Classifier Eval performs significantly better than the Current Eval. This further demonstrates that the misalignment issue is highly manageable. Such a lightweight and computationally feasible post-processing operation facilitates the evaluation of $f_{m}$.
@@ 6 Experiments
6.1 Datasets
@@ 6.2 Implementation Details
For the backbone of models, $f_{b}$, we primarily use bert-base-uncased Devlin et al. (2018) as the base architecture, and the fine-tuned checkpoints are referenced from Yang et al. (2024b). Additionally, for a more comprehensive evaluation, we also adopt T5-base and T5-large Raffel et al. (2020) as $f_{b}$. For the scaling hyperparameters of the Task Arithmetic and Ties models, we follow the original paper by using a portion of labeled data for selection. These parameters are then fixed across all evaluation protocols to ensure a fair comparison. Our code is released in Github  [^1] https://github.com/fskong/FT-Classifier-for-Model-Merging. For a more comprehensive analysis of the misalignment issue and the effectiveness of FT-Classifier Eval, we also introduce experiments on CV tasks, shown in Appendix A.2.
6.3 Results Analysis
@@
Each merging model is evaluated on all datasets using three protocols: Current Eval, KNN Eval and FT-Classifier Eval. To save space, we compute the average accuracy and present it using a line chart. To demonstrate that the FT-Classifier Eval maintains a certain level of robustness even with a small number of few-shot samples, we use five random seeds and report the mean and variance.
The experimental results using BERT as $f_{b}$ have been reported in Figure 2, Figure 5 and Table 2. Unless otherwise specified, the experiments in this paper are based on BERT. All these results prove the existence of misalignment, and FT-Classifier Eval could alleviate this issue, achieving a better assessment of merging methods.
@@
Results conducted on T5-base and T5-large are shown in Figure 6. From these results, we find that the main conclusions are essentially consistent with those on BERT, strengthening our arguments.
6.4 $f_{m}$ vs. $f_{b}$
@@
Previous analyses demonstrate that different $f_{m}$ achieves significant improvements when evaluated with an aligned classifier. To further verify that the performance improvement stems from the inherent capability of the $f_{m}$ itself rather than additional information introduced by fine-tuning, we tested the results of the original pre-trained models $f_{b}$ under different evaluation protocols. The experimental results are presented in Table 3.
@@
From the Table, we observe that regardless of the evaluation protocol, $f_{m}$ consistently outperforms $f_{b}$. This demonstrates that all $f_{m}$ models effectively incorporate knowledge from fine-tuned models. Meanwhile, FT-Classifer Eval does not show significant improvements over Current Eval when applied to $f_{b}$, indicating that models with limited parameters (only the classifier) cannot retain sufficient task-specific information. This proves that FT-Classifer Eval just activates the fine-tuned information that $f_{m}$ already contains.
@@
6.5 FT-Classifier vs. Surgery
The related work Surgery Yang et al. (2024b) characterizes the overall difference between the merging outputs and the fine-tuned outputs as “representation bias”, and reduces this bias by training additional parameters on a large amount of unlabeled data.
We argue that merging models naturally alters the output distribution (representation bias) but does not necessarily harm classification performance. The main cause of performance decline is the misalignment between the merging output and the fine-tuned classifier. We also show that this misalignment is simple (converging to an orthogonal transformation) and does not affect the intrinsic properties of merging outputs. Consequently, it can be mitigated by fine-tuning an aligned classifier for each task with very few samples. In addition, the FT-Classifier can also keep the parameter scale and network structure unchanged. In the context of model merging, we believe that such a lightweight and computationally efficient approach is more favorable. Results in Figure 7 demonstrate that the FT-Classifier outperforms Surgery when only a small number of samples are available.
@@ 7 Related Works
The task vector indicates the offset of a certain task based on the pre-trained model in the model parameter space. We introduce the following latest methods except for the representative referred to before. Motivated by model pruning that eliminating unnecessary parameters could also maintain performance, Dare Yu et al. (2024) reduces potential conflicts in the delta parameters by randomly dropping. To mitigate forgetting while training, MagMax Marczak et al. (2024) combines sequential fine-tuning with maximum-magnitude weight selection to enable effective knowledge integration across tasks. TALL-masks Wang et al. (2024) localizes task-specific information within the multi-task vector by deactivating irrelevant parts for each task using binary masks. Model Tailor Zhu et al. (2024) derives a sparse mask to identify the “model patch” using a fusion of salience and sensitivity analysis, then introduces a compensation mechanism to “decorate the patch”, enhancing performance on both target and original tasks.
@@ 8 Conclusion
Model merging combines multiple fine-tuned models into a single one in the parameter space, enabling multi-task ability without full datasets or extra fine-tuning. In this paper, we first identify the misalignment issue between merging outputs and fine-tuned classifiers in classification tasks and show that merging outputs already contain necessary classification information. We then prove that this misalignment can be converged as an orthogonal transformation, offering insights into misalignment alleviation. We finally introduce an enhanced evaluation protocol, FT-Classifier Eval, by fine-tuning an aligned classifier with few-shot examples, boosting evaluation quality with minimal time cost.
@@ A.1 Visualization of Sentence Embedding
Refer to caption Refer to caption
(a) CV tasks performance on ViT-B-16 (b) CV tasks performance on ViT-B-32
Figure 11: The performance of different evaluation protocols on CV tasks under ViT-B-16 and ViT-B-32.
A.2 Comparison on CV Tasks
A.2.1 Datasets
A.2.2 Implementation Details
For $f_{b}$, we primarily utilize ViT-B-16 and ViT-B-32 Radford et al. (2021) as the base architecture, and the fine-tuned checkpoints are referenced from Yang et al. (2024b). The selection for the merging scale is consistent with the description in the NLP tasks. It is worth noting that we found that, possibly due to the large number of classes in most CV datasets, using few-shot samples is insufficient to fine-tune the classifier effectively. Therefore, we increased the number of samples used for misalignment alleviation. Specifically, we sampled subsets from the test set with sampling rates of {0.5%, 1%, 2.5%, 5%, 10%}.
In the ViT series of experiments, the image input is processed through the image encoder, and text inputs (label names) are processed through the text encoder, obtaining the image embedding $\mathbf{E}_{image}$ and text embeddings $\mathbf{E}_{text}$ respectively. Classification is performed by matching $\mathbf{E}_{image}$ with $\mathbf{E}_{text}$. In the settings of most model merging papers, the text encoder is frozen, meaning that $\mathbf{E}_{text}$ remains unchanged. In this case, we consider $\mathbf{E}_{text}$ can be understood as the classifier parameter $\mathbf{W}_{t}$, that is, initializing the classifier with label name embeddings. This approach is consistent with the way many text classification methods initialize label embeddings.
A.2.3 Visualization of Image Embedding
The image embeddings produced by the $f_{m}$ model are shown in Figure 9 and 10. It is worth noting that, due to the large number of classes in the CV datasets, we did not display the T-SNE results of the fine-tuned model for better visualization. As shown in these Figures, we find that the embeddings produced by different $f_{m}$ also exhibit good clustering performance, which is consistent with the conclusions drawn from the NLP tasks. Therefore, we believe that the misalignment problem also exists in CV tasks.
A.2.4 Results Analysis
Further, to mitigate the misalignment issue in CV tasks, we applied the KNN Eval and FT-Classifier Eval to re-evaluate different merging methods. The comparable results are shown in Figure 11. We find that with just over 1% of unlabeled data, the misalignment problem can be alleviated. This not only demonstrates the effectiveness of the FT-Classifier Eval but also implicitly proves that misalignment is easy to capture.